OpenStudy (anonymous):

If h(x) = 2x + 1, find the value of x if h(x) = 11

OpenStudy (anonymous):

11=2x+1 -1 -1 10=2x divide both sides by 2 x=5

OpenStudy (anonymous):

h(x)=2x+1 -eq. 1 h(x)=11 -eq. 2 from eq. 1 and 2 2x+1=11 2x=10 x=10/2=5

OpenStudy (anonymous):

x=5 :)